OpenStudy (anonymous):

Find the Volume of a triangular prism with a height of 5cm and a base that is a regular triangle with sides of 3cm.

OpenStudy (anonymous):

@85295james

OpenStudy (anonymous):

Yo come to this one @85295james

OpenStudy (anonymous):

The volume of a rectangular prism can be found by the formula: volume=length*width*height

OpenStudy (anonymous):

how do i plug it in?

OpenStudy (anonymous):

@85295james

OpenStudy (anonymous):

the area of any triangle is 1/2bxh. The area of a triangular prism is this x 5cm. Using the pythagorean theorem we know that (1/2(3))^2 + h^2 = 3^2 2.25 -9 = -h^2 6.75 = h^2 (6.75)^1/2 = h = 2.598 Therefore the volume = (1/2)x(3)x(2.598)x(5) = 19.49 cubic cm.

OpenStudy (anonymous):

No decimals xD ill write it down